OpenStudy (anonymous):

Increasing life expectancy. A smoker is 15 times more likely to die of lung cancer than a nonsmoker. An ex-smoker who stopped smoking t years ago is w times more likely to die of lung cancer than a nonsmoker, where w = 15 – t. Graph the equation and use the graph to estimate how much more likely it is for Sandy to die of lung cancer than Polly, if Polly never smoked and Sandy quit 2 1/2 yrs ago.

OpenStudy (anonymous):

polly is nonsmoker sandy is ex- smoker t= 2 1/2= 5/2 years w= 15-t = 15- 5/2 = 25/2 so sandy is 25/2 more times likely to die than polly

OpenStudy (anonymous):

so how do i graph this answer

OpenStudy (anonymous):

im so lost

OpenStudy (anonymous):

do you wanna draw graph,,?

OpenStudy (anonymous):

yes, what points to draft?

OpenStudy (anonymous):

take t on x-axis starts from 0 to ....at origin and w on y-axis starts from 0 to...15 now match the respective points of sandy and polly

OpenStudy (anonymous):

so plot (0,0)?

OpenStudy (anonymous):

thats origin...for polly..

OpenStudy (anonymous):

so what nmber is the plot for sandy?

OpenStudy (anonymous):

sorry ...t=0 so w= 15-0=15 for polly so (0,15) plolly and (5/2,25/2) sandy

OpenStudy (anonymous):

the graph will be up to down starting from 15 on y - axis

OpenStudy (anonymous):

i know how to plot for y

OpenStudy (anonymous):

but x how do i plot 5/2 and 25/2, when i divide it comes out to decimals, it doesn't make sense to me...

OpenStudy (anonymous):

yes ..it will come in decimal. dont you know how to pint the decimal..?

OpenStudy (anonymous):

*point

OpenStudy (anonymous):

no i don't....

OpenStudy (anonymous):

are you kidding or seriously?..lol

OpenStudy (anonymous):

oh nevermind lol...

OpenStudy (anonymous):

math is not my favorite so i tend to forget certain things, but thanks hon you are a sweetie

OpenStudy (anonymous):

my pleasure....it was my very first day here ...nice to meet you..:P

OpenStudy (anonymous):

same here i may need you again lol, you so freaken smart....

OpenStudy (anonymous):

okay. you can mail me here mr.mukesh009@gmail.com. can i know your name if you dont mind..?

OpenStudy (anonymous):

Ebony

OpenStudy (anonymous):

do you have a FB page?

OpenStudy (anonymous):

yes. you can join me in fb through the link below http://www.facebook.com/profile.php?id=100000926381100